# RDDM: Practicing RAW Domain Diffusion Model for Real-world Image Restoration

### v1

@@ Abstract
We present the RAW domain diffusion model (RDDM), an end-to-end diffusion model that restores photo-realistic images directly from the sensor RAW data. While recent sRGB-domain diffusion methods achieve impressive results, they are caught in a dilemma between high fidelity and realistic generation. As these models process lossy sRGB inputs and neglect the accessibility of the sensor RAW images in many scenarios, e.g., in image and video capturing in edge devices, resulting in sub-optimal performance. RDDM bypasses this limitation by directly restoring images in the RAW domain, replacing the conventional two-stage image signal processing (ISP)IR pipeline. However, a simple adaptation of pre-trained diffusion models to the RAW domain confronts the out-of-distribution (OOD) issues. To this end, we propose: (1) a RAW-domain VAE (RVAE) learning optimal latent representations, (2) a differentiable Post Tone Processing (PTP) module enabling joint RAW and sRGB space optimization. To compensate for the deficiency in the dataset, we develop a scalable degradation pipeline synthesizing RAW LQ-HQ pairs from existing sRGB datasets for large-scale training. Furthermore, we devise a configurable multi-bayer (CMB) LoRA module handling diverse RAW patterns such as RGGB, BGGR, etc. Extensive experiments demonstrate RDDM’s superiority over state-of-the-art sRGB diffusion methods, yielding higher fidelity results with fewer artifacts.
@@ 1 Introduction
Real-world Image Restoration (Real-IR) aims to restore high-quality (HQ) images from low-quality (LQ) images containing complex degradations, e.g. noise, image compression and blur (Fan et al., 2020; Jinjin et al., 2020; Zhang et al., 2022; 2019; 2018b; 2017). Existing GAN-based (Ledig et al., 2017; Wang et al., 2018) methods employ a generator and a discriminator for adversarial training. However, GAN-based methods suffer from pattern collapse, incurring unsatisfactory results (Wang et al., 2021; Zhang et al., 2021; Liang et al., 2022a; Chen et al., 2022; Liang et al., 2022b; Xie et al., 2023). Benefiting from the powerful generative priors granted by text-to-image (T2I) models, SUPIR (Yu et al., 2024) and its counterparts (Zhang et al., 2023; Lin et al., 2024; Wang et al., 2024; Wu et al., 2024b; Yu et al., 2024; Sun et al., 2024; Menon et al., 2020; Karras et al., 2019) integrate T2I models into Real-IR and have attained remarkable performance. Nevertheless, the dilemma between image fidelity and realistic generation remains a pivotal challenge that these methods must confront.
Notably, prevailing Real-IR models process and enhance images in the lossy sRGB domain. This may lead to sub-optimal results in many scenarios where the sensor RAW images are accessible, e.g., in image and video capturing in edge devices, as the rich meta-info contained in the sensor RAW images is not effectively squeezed. This drives us to integrate the powerful pre-trained T2I models with sensor RAW data.
However, a naive combination of prior models trained in the sRGB domain with the sensor RAW images confronts substantial challenges. Firstly, sRGB images and sensor RAW images differ significantly in terms of luminance, mosaic patterns of RAW images and noise distribution, as shown in Fig. 1, so that publicly available VAE pre-trained in the sRGB domain cannot effectively encode and decode sensor RAW images. Secondly, a comprehensive RAW domain Real-IR dataset that could serve as a rigorous benchmark for training and evaluating IR models in the RAW domain is still absent.
To address these issues, in this paper we propose the RAW domain diffusion model (RDDM). As a remedy to the first challenge, we devise a RAW domain VAE (RVAE) that accepts a sensor RAW as the input and outputs a clean image in the linear domain. We employ a divide-and-conquer strategy to train the RVAE. Initially, we fine-tune a pre-trained VAE using a linear HQ dataset to bridge the gap between the sRGB and linear domains. Subsequently, we adopt the LoRA fine-tuning approach to jointly train the encoder of the RVAE and Real-IR, thereby adapting it to the mosaic pattern of real sensor RAW data. To further mitigate out-of-distribution (OOD) issues, we design a differentiable post tone processing (PTP) module that enables joint RAW and sRGB space optimization, thereby endowing the model with improved fidelity. Additionally, a configurable multi-Bayer (CMB) LoRA module is designed to adapt our model to different Bayer patterns of RAW. To compensate for the deficiency in the dataset, we propose a realistic RAW domain degradation pipeline that amalgamates the degradation strategies of ESRGAN (Wang et al., 2021) and UPI (Brooks et al., 2019), allowing us to synthesize abundant RAW–linear image pairs from publicly accessible sRGB datasets.
In summary, our main contributions are as follows:
@@
We propose RDDM, the first practical application of the raw domain diffusion model, establishing a novel paradigm for RAW image restoration.
We propose RVAE capable of encoding mosaiced RAW images and subsequently decoding the latent representations into linear HQ images in order to address the OOD issues.
Additionally, we design a RAW domain Real-IR data synthesis pipeline and construct a RAW Real-IR benchmark.
Extensive experiments verify that RDDM demonstrates superior image fidelity and comparable generation capability to the state-of-the-art methods.
@@ 2 Related Work
Real-world Image Restoration. Real-IR is becoming a trending field of research since the advent of ESRGAN (Ledig et al., 2017). Early studies attempted various ways to combine generative adversarial networks (GANs) (Goodfellow et al., 2014; Karras et al., 2017; 2019; Radford et al., 2015; Mirza & Osindero, 2014) with perceptual losses (Ding et al., 2020; Johnson et al., 2016; Zhang et al., 2018a) for training networks to predict images that follow the natural image distribution (Ledig et al., 2017; Wang et al., 2018; 2021; Zhang et al., 2021; Liang et al., 2022a; Chen et al., 2022; Liang et al., 2022b; Xie et al., 2023). However, since adversarial training of GANs can be unstable, their discriminators are deficient in determining the quality of the diverse natural image contents, giving rise to unnatural visual artifacts. As an alternative to GAN-based methods, diffusion-based models (Podell et al., 2023; Rombach et al., 2022) is becoming increasingly popular in Real-IR tasks (Kawar et al., 2022; Li et al., 2022; Luo et al., 2023a; b; Özdenizci & Legenstein, 2023; Saharia et al., 2022) to generate realistic images with substantial texture, leveraging pre-trained Stable Diffusion (SD) models as priors (Zhang et al., 2023; Lin et al., 2024; Wang et al., 2024; Wu et al., 2024b; Yu et al., 2024; Sun et al., 2024; Menon et al., 2020; Karras et al., 2019) whereas they employ different condition injection strategies and feature extraction. Nevertheless, all existing Real-IR methods restore images in the sRGB domain in which rich information in the RAW domain might be lost after ISP. However, directly adapting sRGB Real-IR methods to the RAW domain encounters severe domain mismatch and results in poor performance.
Sensor RAW Images. Sensor RAW refers to the unprocessed, original data collected directly from specific camera sensors, holding a wealth of physical information. Sensor RAW contains noise and the Bayer Color Filter Array (CFA) patterns (Snyder et al., 1995; Beenakker & Patra, 1999; Gotoh & Okutomi, 2004; Maschal Jr et al., 2010). The demosaicing process removes the Bayer CFA pattern and produces linear domain images from RAW, and the Post Tone Processing (PTP) module produces natural sRGB image from linear domain images. In short, sensor RAW images differ from sRGB ones in terms of sensor-captured information (e.g. 12-16 bit photon-electric signals), noise distribution, color space, luminance, dynamic range and image format (e.g. mosaic patterns), making it difficult for existing sRGB Real-IR methods to adapt to the RAW domain.
Image Signal Processing. An ISP pipeline reconstructs a visually appealing sRGB image from RAW sensor data. Traditionally, an ISP pipeline is formulated as a series of hand-crafted modules executed sequentially (Sundararajan, 2017), including some representative steps such as demosaicing (DM), denoising (DN), automated-white-balance (AWB), color correction matrix (CCM), gamma compression (GC), and tone mapping (TM). Among these steps, DM (Li et al., 2008; Alleysson et al., 2005; Hirakawa & Parks, 2005; Kimmel, 1999) and DN modules (Brooks et al., 2019; Cao et al., 2024; Li et al., 2024) are ill-posed and can cause over-smoothen images (Qian et al., 2019) and AWB, CCM and TM modules can cause information compression due to their inherent data clipping operations. In addition, different digital imaging device producers adopt different ISP pipelines which usually remain as black boxes, making it difficult to obtain information about the specific steps inside.
@@ 3 Methodology
3.1 Preliminaries
Problem Modeling. Real-world image restoration in the sRGB domain aims to train a neural network , parameterized by , estimating HQ sRGB image given LQ sRGB image . In RAW domain, we train a neural network to transform LQ sensor RAW to HQ linear domain image . The training task can be modeled as the following optimization problem:
@@ 3.1 Preliminaries
Image Signal Proccessing. Given a sensor RAW with single channel as input, DN and DM module produces linear domain image with three channels, which can be formulated as:
where is the inverse function, transfers the linear domain image to the RAW domain image . Post Tone Processing (PTP) module including AWB, CCM, GC and TM, converts linear domain images to sRGB images , which is defined as:
where strips the color information from an sRGB-encoded image and converts the resulting grayscale representation into the linear domain. A detailed introduction of ISP and Inverse ISP, as well as their mathematical derivations referenced in the Appendix.
@@ 3.2 RAW Domain Diffusion Model
Framework Overview. Our generator is composed of an RVAE encoder , a diffusion network , and an RVAE decoder . extracts the latent features of the sensor RAW image. jointly optimizes DN, DM, and image restoration in the latent space to obtain the latent features. then decodes the latent features to produce the linear domain image, which is subsequently mapped to the sRGB domain by the PTP module. We incorporate trainable LoRA layers (Hu et al., 2022) into the pre-trained and . To address the issue of extracting prompts from sensor RAW images, the feed-forward ISP firstly processes the sensor RAW image to obtain an sRGB image, from which the DAPE (Zheng et al., 2024) prompt extractor extracts the textual information to activate priors in the model training. Fig. 3 presents an overview of the framework and the interplay between the various modules.
RAW Domain VAE. VAE plays a pivotal role in the quality of generated images. However, existing VAE in the sRGB domain are not capable of effectively encoding RAW images and decoding linear domain images. Therefore, we train a RAW domain VAE that encodes RAW images and subsequently decodes the latent representation into a linear domain image. (Rombach et al., 2022) employs a scaling factor to normalize the latent space distributions of different VAEs to the standard Gaussian distribution, which is beneficial for diffusion network optimization. To obtain the statistically accurate scaling factor, we calculate the parameter for training samples in the linear domain according to the following formula:
where denotes the latent space of the training samples encoded by . and present the mean and variance of the data distribution. The rescaled latent has unit standard deviation, i.e., .
The training strategy of RVAE is illustrated in Fig. 3 (a). We train the encoder and decoder on the linear domain dataset, such that the linear domain input is encoded by RVAE encoder to obtain the latent feature and the RVAE decoder decodes into the target linear domain image . Furthermore, we introduce a differentiable PTP module that simultaneously supervises training in both the sRGB and RAW domains. Similar to LDM (Rombach et al., 2022), we use loss, LPIPS loss, and GAN loss to train the VAE encoder and decoder to generate realistic details of a linear image:
where and is calculated in both the RAW and sRGB domains. and represents the gradient of the last layer in the decoder.
Training Framework. In order to accommodate RAW images captured with arbitrary Bayer patterns, we propose a configurable multi-Bayer (CMB) LoRA module that augments the RVAE encoder and the pre-trained diffusion network with independent sets of LoRA, and we assign a distinct LoRA group to each Bayer pattern. During training, the RVAE decoder is frozen and only the CMB modules are optimized. Consequently, the encoder learns to encode RAW images and diffusion network jointly performing DN, DM, and detail enhancement. We use VSD loss, LPIPS loss, and MSE loss to train our model in the RAW domain and the sRGB domain:
@@
Feed-forward ISP. To supervise the training of RDDM in sRGB domain and extract prompt information from RAW images, we devise a feed-forward ISP that can transform RAW images into the sRGB domain, as illustrated in Fig. 4 (b). The required metadata, including Bayer pattern, AWB gain, and CCM, are obtained from Inverse ISP and subsequently employed as parameters of the feed-forward ISP. We train a lightweight neural network DDNet to replace conventional DN and DM modules by minimizing: , where is the joint DN and DM network, is the sensor RAW and is the HQ linear image.
RAW prompt extractor. To address the issue that existing text extractors fail to accurately extract text information from sensor RAW images with significant noise, we propose the RAW prompt extractor (RPE). Initially, we train a lightweight denoising and demosaicing network to transform noisy sensor RAW images into clean linear domain ones. Subsequently, the PTP module converts these linear domain images into sRGB images. Finally, we employ the DAPE to accurately extract the prompt information. can significantly reduce the impact of noise on text extractors, thereby enhancing the robustness of the prompt extractor model.
3.3 Image Synthesis Pipeline
Synthetic Image Degradation Pipeline. Despite the abundance of existing datasets for Real-IR, such as LSDIR (Li et al., 2023), FFHQ (Karras et al., 2019), and DIV2K (Agustsson & Timofte, 2017), these datasets are all in the sRGB domain. To the best of our knowledge, there is currently no dataset for Real-IR in the RAW domain. Therefore, to provide a solid training foundation for Real-IR in the RAW domain, we synthesize a RAW domain Real-IR dataset by degrading publicly available sRGB Real-IR datasets through our synthetic image degradation pipeline, as shown in Fig. 4 (a). In particular, we first degrade sRGB HQ images to sRGB LQ images via detail degradation method, following Real-ESRGAN (Wang et al., 2021) despite excluding the degradation process of random noise, since the noise in the RAW domain is intrinsic in the sensor’s physical features. Subsequently, for the synthesis of the training dataset for RDDM, the sRGB LQ images are processed through the inverse post tone processing module (IPTP) and mosaic noise synthesizer module (MNS) to obtain degraed RAW images , while the sRGB HQ images are transformed into linear domain GT through inverse PTP module. For the synthesis of the training dataset for DDNet, linear HQ images are processed through MNS module to produce detailed RAW images .
@@ 4.1 Experimental Settings
Training and Testing Datasets. We adopt the OSEDiff (Wu et al., 2024a) setup and train RDDM using the LSDIR (Li et al., 2023) dataset and the first 10K face images from FFHQ (Karras et al., 2019). We use the degradation pipeline discussed in the Image Synthesis Pipeline section to synthesize LQ and HQ pairs in the RAW domain. For testing, we construct our benchmark by degrading the HR images from the DIV2K-Val, consisting of 100 images, RealSR containing 100 images, and DRealSR containing 93 images, using our proposed data synthesis pipeline.

### v2

@@ Abstract
We present the RAW domain diffusion model (RDDM), an end-to-end diffusion model that restores photo-realistic images directly from the sensor RAW data. While recent sRGB-domain diffusion methods achieve impressive results, they are caught in a dilemma between high fidelity and image generation. As these models process lossy sRGB inputs and neglect the accessibility of the sensor RAW images in many scenarios, e.g., in image and video capturing in edge devices, resulting in sub-optimal performance. RDDM obviates this limitation by directly restoring images in the RAW domain, replacing the conventional two-stage image signal processing (ISP)Image Restoration (IR) pipeline. However, a simple adaptation of pre-trained diffusion models to the RAW domain confronts many challenges. To this end, we propose: (1) a RAW-domain VAE (RVAE), encoding sensor RAW and decoding it into an enhanced linear domain image, to solve the out-of-distribution (OOD) issues between the different domain distributions; (2) a configurable multi-bayer (CMB) LoRA module, adapting diverse RAW Bayer patterns such as RGGB, BGGR, etc. To compensate for the deficiency in the dataset, we develop a scalable data synthesis pipeline synthesizing RAW LQ-HQ pairs from existing sRGB datasets for large-scale training. Extensive experiments demonstrate RDDM’s superiority over state-of-the-art sRGB diffusion methods, yielding higher fidelity results with fewer artifacts. Codes will be publicly available at github.com/YanCHEN-fr/RDDM.
@@ 1 Introduction
Real-world Image Restoration (Real-IR) aims to restore high-quality (HQ) images from low-quality (LQ) images containing complex degradations Fan et al. (2020); Zhang et al. (2022); Jinjin et al. (2020); Zhang et al. (2019; 2018b; 2017). Real-IR has transitioned from GAN-based adversarial training Ledig et al. (2017); Wang et al. (2018); Xie et al. (2023) to leveraging the powerful generative priors of text-to-image (T2I) diffusion models Zhang et al. (2023); Lin et al. (2024); Wang et al. (2024); Wu et al. (2024b); Yu et al. (2024). Despite their success in the sRGB domain, these methods encounter a fundamental bottleneck: the Image Signal Processor (ISP) irreversibly discards critical details and compresses the dynamic range from high-bit (12-14 bits) RAW images into 8-bit sRGB images. Restoring HQ images from such degraded sRGB inputs inevitably leads to sub-optimal results Nguyen and Brown (2016); Xing et al. (2021).
To transcend the performance ceiling of Real-IR, shifting restoration to the sensor RAW domain, i.e. the front-end of ISP, holds substantial promises. RAW data preserves the original optical and physical signals, offering higher bit depth and richer information. Nevertheless, integrating T2I models into the RAW domain is confronted with two critical challenges: (1) Distribution Disparity: the significant gaps in the luminance and mosaic patterns between the sRGB and RAW images, as shown in Fig.3, make direct application of T2I models in the RAW domain ineffective. (2) Generalization Constraints: varying sensor Bayer patterns and noise characteristics hinder the ability of the model to generalize across different devices.
@@
In this paper, we propose the RAW Domain Diffusion Model (RDDM), the first practical diffusion-based paradigm for RAW image restoration. To address the aforementioned challenges, we introduce three core innovations:
RVAE: we design a RAW domain Variational Autoencoder (RVAE) using a divide-and-conquer strategy, mapping noisy mosaicked RAW inputs into a latent space and decoding them into enhanced linear images to bridge the RAW-sRGB distribution gap.
CMB-LoRA: a sensor-aware adaptation module that strategically selects LoRA weights to handle heterogeneous Bayer patterns, significantly enhancing cross-sensor generalization.
RAW Image Synthesis Pipeline: to overcome data scarcity, we propose a synthesis method to generate high-fidelity RAW-linear pairs from large-scale sRGB datasets.
Extensive experiments demonstrate that RDDM achieves state-of-the-art fidelity and generative performance, as illustrated in Fig. 3.
@@ 2 Related Work
Real-world Image Restoration. Since the advent of ESRGAN Ledig et al. (2017), Real-IR has evolved from GAN-based approaches Ledig et al. (2017); Wang et al. (2018; 2021); Zhang et al. (2021); Liang et al. (2022); Chen et al. (2022), which often suffer from training instability and unnatural artifacts, to diffusion-based models. By leveraging pre-trained Stable Diffusion (SD) priors, the latter achieve superior visual realism and texture generation through various condition injection strategies Podell et al. (2023); Rombach et al. (2022); Wang et al. (2024); Wu et al. (2024b); Yu et al. (2024). However, existing Real-IR methods perform predominantly in the sRGB domain, where critical sensor-level information is often discarded by the ISP. Furthermore, directly transferring sRGB-trained models to the RAW domain yields poor performance due to severe domain mismatch. This gap necessitates a specialized framework that effectively adapts the generative power of diffusion models to the unique characteristics of RAW data.
Image Processing in RAW Domain. Modern imaging sensors capture RAW data, which is converted to sRGB by a traditional ISP pipeline via sequential modules, including demosaicing (DM), denoising (DN), and tone mapping (TM). Subsequent post-processing (e.g., AWB, clipping) compresses dynamic range, leading to irreversible detail loss in the final sRGB output. One-stage methods Brooks et al. (2019) typically integrate DN and DM but often suffer from over-smoothing due to limited representation capacity. In contrast, integrating diffusion models into the RAW domain harnesses essential sensor signals and enhances generative capacity, providing a superior foundation for high-fidelity image restoration.
@@ 3 Methodology
3.1 Problem Modeling
In the sRGB domain, Real-IR model , parameterized by , aims to estimate HQ sRGB image given LQ sRGB image . In RAW domain, we train a neural network to transform LQ sensor RAW to HQ linear domain image . The training task can be modeled as the following optimization problem:
@@ 3.1 Problem Modeling
Existing AIGC-based image restoration paradigms typically follow the pipeline illustrated in Fig. 4 (a). The initial sensor RAW is processed through conventional image signal processing, which integrates a denoising, demosaicing module (DDM), and post-tone mapping modules to generate an sRGB image. The AIGC model is subsequently applied to this sRGB output to perform high-quality enhancement. Formally, this paradigm can be expressed as:
In contrast, our proposed RDDM approach, depicted in Fig. 4 (b), bypasses the traditional denoising and demosaicing modules and shifts the AIGC-driven enhancement process into the RAW domain. Following the enhancement, the output of the AIGC model is transformed into the sRGB domain via a post tone mapping (PTP) module. Formally, the RDDM paradigm can be defined as:
By operating directly on RAW data, RDDM mitigates the irreversible information loss inherent in standard ISP transformations. Since on-device ISP pipelines are typically proprietary ”black boxes,” a comprehensive overview of ISP components and their corresponding mathematical derivations is provided in Appendix A.1.
@@ 3.2 RAW Domain Diffusion Model
Framework Overview. Our framework integrates an RVAE and a diffusion network for high-fidelity restoration. Specifically, the encoder maps mosaicked RAW inputs into a latent space, where performs iterative refinement to recover fine-grained details. The decoder then reconstructs these features into the linear domain, followed by an sRGB transformation via the PTP module. To handle diverse Bayer configurations, we incorporate trainable LoRA layers Hu et al. (2022) into the pre-trained and , facilitating adaptive adaptation through our CMB-LoRA strategy. Furthermore, to provide semantic guidance, we employ a DAPE Zheng et al. (2024) to derive textual prompts from sRGB images rendered by the ISP, thereby activating generative priors. An overview of the framework is depicted in Fig. 5.
RAW Domain VAE. Owing to the vast domain shift, existing sRGB domain VAEs fail to reconstruct effectively, resulting in diffusion model-generated images characterized by low visual quality and significant color shifts. Therefore, we devise a RAW domain VAE that encodes RAW images and subsequently decodes the latent representation into a linear domain image. The training of the RVAE is conducted in two distinct stages. To ensure the decoder is capable of reconstructing images in the linear domain, the first stage involves training on a reconstruction task utilizing linear domain images, following the strategy illustrated in Fig. 5 (a). Given a linear domain input , the encoder extracts the latent representation , which is then reconstructed by the decoder as . Furthermore, we devise a differentiable PTP module that simultaneously supervises training in both the sRGB and RAW domains. Similar to LDM Rombach et al. (2022), we use loss, LPIPS loss, and GAN loss to train the VAE encoder and decoder to generate realistic details of a linear image:
where and is calculated in both the RAW and sRGB domains. and represents the gradient of the last layer in the decoder. Following Rombach et al. (2022), to facilitate diffusion model optimization, we normalize the latent distribution to a standard Gaussian. The scaling factor is computed over the training samples:
where denotes the latent space of the training samples encoded by . and present the mean and variance of the data distribution. The rescaled latent has unit standard deviation, i.e., . To enable the RVAE to process RAW images, the second training stage involves freezing the encoder and decoder weights optimized in the first stage and incorporating LoRA into the encoder. Concurrently, LoRA modules are integrated into the diffusion model to facilitate the removal of mosaic artifacts and heavy noise while simultaneously enhancing fine-grained image details. The overall training strategy is illustrated in Fig. 5 (b). We use VSD loss, LPIPS loss, and MSE loss to train our model in the RAW domain and the sRGB domain:
@@
CMB-LoRA. To tackle the traditional cross-device generalization challenge posed by diverse Bayer patterns and noise characteristics, we propose Configurable Multi-Bayer (CMB) LoRA. Rather than costly full-parameter retraining, our approach incorporates pattern-specific LoRA branches into the RVAE and diffusion networks. Each branch is independently optimized for a specific Bayer configuration during training and adaptively selected during inference. This parameter-efficient strategy allows our model to handle diverse sensor characteristics without retraining from scratch, as detailed in Fig. 5 (b).
3.3 Data Synthesis
Despite the prevalence of sRGB datasets for Real-IR (e.g., LSDIR Li et al. (2023), DIV2K Agustsson and Timofte (2017)), a large-scale RAW domain dataset remains absent. To address this deficiency, we construct a large-scale synthetic RAW-domain dataset by applying a degradation pipeline to publicly available sRGB resources, thereby facilitating effective training in the RAW domain, as shown in Fig. 6. Following Real-ESRGAN Wang et al. (2021), we first apply detail degradation to obtain from . Crucially, we exclude sRGB-specific JPEG compression and random noise, as these do not align with native sensor characteristics. For stage one training, an inverse post tone processing (IPTP) module projects into the linear domain to obtain the RVAE target . In stage two, we synthesize RAW domain LQ inputs by passing through the IPTP module and a mosaic noise synthesizer (MNS). Throughout training, serves as the ground truth. The process is formulated as:
@@ 4.1 Experimental Settings
Training and Testing Datasets. RDDM is trained on LSDIR Li et al. (2023) and a 10K-image subset of FFHQ Karras et al. (2019). The RVAE is trained on a broader collection including DIV2K Timofte et al. (2017), Flickr2K Agustsson and Timofte (2017), LSDIR Li et al. (2023), DIV8K Gu et al. (2019), and FFHQ Karras et al. (2019). RAW-domain training pairs are synthesized following the pipeline in Sec. 3.3. For evaluation, we employ both real-world and synthetic benchmarks. Real-world RAW testing is conducted on DND Plotz and Roth (2017), RealCapture Li et al. (2024), and SIDD Abdelhamed et al. (2018), covering diverse consumer cameras and mobile devices. Additionally, synthetic RAW benchmarks are constructed by applying our degradation method to HR images from DIV2K-Val Timofte et al. (2017), RealSR Cai et al. (2019), and DRealSR Wei et al. (2020).
Compared Methods. We compare RDDM against the best-performing traditional one-stage method and two-stage ISPIR models methods, as shown in Table 1. For one-stage methods, we re-train JDnDmSR (Xing and Egiazarian, ), SwinIR (Liang et al., 2021), and MambaIRv2 (Guo et al., 2025) on the same RAW domain training dataset as our baseline. For two-stage ISPIR methods, we choose PIPNet (A Sharif et al., 2021) as the DN and DM module for ISP and diffusion-based IR methods as our Real-IR baselines, including StableSR (Wang et al., 2024), DiffBIR (Lin et al., 2024), PASD (Yang et al., 2024), SeeSR (Wu et al., 2024b), SUPIR (Yu et al., 2024), and OSEDiff (Wu et al., 2024a). Additionally, we conduct a comparative analysis with GAN-based Real-IR methods, including BSRGAN Zhang et al. (2021), Real-ESRGAN Wang et al. (2021), LDL Liang et al. (2022), and FeMaSR Chen et al. (2022). The comparisons of RDDM against GAN-based Real-IR may be found in Appendix A.2
Implementation Details. We train RDDM with the AdamW optimizer at a learning rate of . The entire training process spans 150000 steps with a batch size of 16. The rank of LoRA in the RVAE Encoder and the diffusion network is set to 4. We employ DAPE as the sRGB domain text prompts extractor. We set , and for all experiments.
4.2 Comparisons with State-of-the-Arts